1. what is the approximate weight of oil in a container 3 feet high and 9 inches in diameter? (oil weighs approximately 55 pounds per cubic foot).
2. find the area of a rectange which measures 12 feet, 3 inches by 3 feet, 9 inches. The correct answer expressed in square feet is?
3.the diameter of a circle is 6 feet, 3 inches. what is the circumference of the circle? NOTE: to convert decimal fractions to inches in this problem, multiply by 12.

Answer by nerdybill(6958)

1. what is the approximate weight of oil in a container 3 feet high and 9 inches in diameter? (oil weighs approximately 55 pounds per cubic foot).
volume of container = (pi)r^2 * height
radius = 1/3(diameter) = 4.5 inches
Convert that into feet: 4.5 / 12 = 0.375 feet
.
volume of container = (pi)r^2 * height
volume of container = (3.14)(0.375)^2 * 3 = 1.3246875 cubic feet
.
"volume" * 55 = 1.3246875 * 55 = 72.86 pounds
.
2. find the area of a rectange which measures 12 feet, 3 inches by 3 feet, 9 inches. The correct answer expressed in square feet is?
.
convert 12 feet 3 inches to feet:
12 + 3/12 = 12 + 1/4 = 48/4 + 1/4 = 49/4 feet
.
convert 3 feet, 9 inches:
3 + 9/12 = 3 + 3/4 = 12/4 + 3/4 = 15/4
.
Area then is:
(49/4)(15/4) = 735/16 = 45.94 square feet
.
3.the diameter of a circle is 6 feet, 3 inches. what is the circumference of the circle? NOTE: to convert decimal fractions to inches in this problem, multiply by 12.
Circumference = (pi)d
where
pi is 3.14
d is diameter
.
Convert 6 feet, 3 inches to feet:
6 + 3/12 = 6 + 1/4 = 24/4 + 1/4 = 25/4 feet
.
Circumference = (pi)d = (3.14)(25/4) = 78.5/4 = 19.625 feet
or in inches:
19.625 * 12 = 235.5 inches